OpenStudy (anonymous):

simplify C(n,1)+C(n,2)+C(n,3)+... +C(n,n-1)+C(n,n) where C(n,n)=n!/(r!(n-r)!)

OpenStudy (anonymous):

It's like the binomial theorem

OpenStudy (anonymous):

C(n,r)=C(n,n-r)

OpenStudy (anonymous):

\[ \Large (a+b)^n = \sum_{k=0}^n\binom{n}{k}a^{n-k}b^k \]

OpenStudy (anonymous):

Let \(a=1, b=1\) and you get: \[ \Large 2^n = \sum_{k=0}^n\binom{n}{k} \]

OpenStudy (anonymous):

\[ \Large 2^n = \sum_{k=0}^n\binom{n}{k} = \sum_{k=1}^n\binom{n}{k}+\binom{n}{1} \]This gives me\[ \Large \sum_{k=1}^n\binom{n}{k} =2^n- \binom{n}{1} \]

OpenStudy (anonymous):

@bernlim Make sense?

OpenStudy (anonymous):

yeah, thanks!